a firm has the following demand function P=60-0.5Q and its total cost are defined by TC=13+Q. find the maximum revenue


1
Expert's answer
2022-04-18T00:05:00-0400

The revenue function is given as follows:


"R(Q) = Q\cdot P(Q) = Q\cdot (60-0.5Q)=-0.5Q^2 +60Q"

In order to find its maximum, let's set its derivative to 0 and solve for "Q^*" :


"R'(Q^*)=-Q^*+60=0\\
Q^*=60"

The value of maximum revenue is obtained by plugging "Q^*" into "R(Q)":


"R(Q^*)=-0.5\cdot 60^2 + 60\cdot 60=1800"

Answer. 1800.